# Singular Value Scaling: Efficient Generative Model Compression via Pruned Weights Refinement

### v1

@@ Abstract
While pruning methods effectively maintain model performance without extra training costs, they often focus solely on preserving crucial connections, overlooking the impact of pruned weights on subsequent fine-tuning or distillation, leading to inefficiencies. Moreover, most compression techniques for generative models have been developed primarily for GANs, tailored to specific architectures like StyleGAN, and research into compressing Diffusion models has just begun. Even more, these methods are often applicable only to GANs or Diffusion models, highlighting the need for approaches that work across both model types. In this paper, we introduce Singular Value Scaling (SVS), a versatile technique for refining pruned weights, applicable to both model types. Our analysis reveals that pruned weights often exhibit dominant singular vectors, hindering fine-tuning efficiency and leading to suboptimal performance compared to random initialization. Our method enhances weight initialization by minimizing the disparities between singular values of pruned weights, thereby improving the fine-tuning process. This approach not only guides the compressed model toward superior solutions but also significantly speeds up fine-tuning. Extensive experiments on StyleGAN2, StyleGAN3 and DDPM demonstrate that SVS improves compression performance across model types without additional training costs. Our code is available at: https://github.com/LAIT-CVLab/Singular˙Value˙Scaling.
@@ Introduction
Generative models like Generative Adversarial Networks (GANs) (Goodfellow et al. 2014) and Diffusion models (Ho, Jain, and Abbeel 2020) have achieved remarkable performance across various computer vision tasks such as image generation (Karras et al. 2020; Ho, Jain, and Abbeel 2020; Wang et al. 2022), image editing (Baykal et al. 2023; Pehlivan, Dalva, and Dundar 2023; Kawar et al. 2023; Zhang et al. 2023), even video generation (Skorokhodov, Tulyakov, and Elhoseiny 2022; Ho et al. 2022; Blattmann et al. 2023) and 3D generation (Chan et al. 2022; Karnewar et al. 2023). The impressive performance of these generative models, however, often comes at the cost of high memory and computational demands, limiting their real-world applicability. To address these issues, several model compression technique for generative models have been proposed (Liu et al. 2021; Xu et al. 2022; Chung et al. 2024; Fang, Ma, and Wang 2023).
Model compression typically involves two steps: 1) pruning to reduce model size while retaining essential pre-trained knowledge, and 2) fine-tuning to restore performance. Among these, effective pruning has received significant attention for its ability to enhance following fine-tuning step. The retained pre-trained knowledge from pruning can enhance fine-tuning, leading to improved performance and faster convergence without additional training costs. While previous pruning methods (Chung et al. 2024; Fang, Ma, and Wang 2023) effectively maintain model performance, they often focus solely on preserving crucial connections in the pre-trained model, overlooking the impact of pruned weights on subsequent process, leading to inefficient fine-tuning or distillation. This issue becomes more severe as the model’s capacity decreases, and in some cases, the pruned weights results in worse performance compared to random initialization; i.e. slow convergence speed and lower performance. Therefore, addressing these factors is essential for achieving more efficient generative models.
@@
Recently, several studies (Liu et al. 2021; Xu et al. 2022; Chung et al. 2024) have been proposed for compressing unconditional GANs, particularly the StyleGAN family (Karras et al. 2020, 2021), which are the state-of-the-art models in the area. CAGAN (Liu et al. 2021) introduced the first framework for pruning pre-trained StyleGAN models and fine-tuning them via pixel-level and feature-level knowledge distillation. Following this, StyleKD (Xu et al. 2022) further enhanced the fine-tuning process by incorporating a specialized relation loss tailored for StyleGAN. In particular, CAGAN introduced a content-aware pruning technique that preserves connections in the pre-trained model crucial for the semantic part of generated images, while StyleKD tackled output discrepancies problem by inheriting only the pre-trained model’s mapping network and randomly initializing the synthesis network with smaller size. More recently, DCP-GAN (Chung et al. 2024) proposed a diversity-aware channel pruning technique, which preserves connections in the synthesis network that contribute to sample diversity, significantly enhancing both the diversity of generated images and the training speed.
@@
Diffusion models have garnered significant attention for their stable training and impressive generative capabilities. However, their performance incurs high computational costs due to iterative sampling. To enhance efficiency, various sampling techniques have been developed to reduce the number of required iterations (Song, Meng, and Ermon 2021; Lu et al. 2022; Zheng et al. 2023). Orthogonal to the these sampling techniques, Diff-Prune (Fang, Ma, and Wang 2023) introduced a seminal work for reducing computational costs by compressing Diffusion models. This demonstrated that by primarily pruning connections involved in less important diffusion steps, a reduced-size Diffusion model with minimal performance loss can be achieved in significantly fewer training iterations compared to the original-sized model.
@@
Weight initialization is essential for efficient training in deep learning. Glorot (Glorot and Bengio 2010) and He initialization (He et al. 2015) are widely used to maintain activation variance and ensure stable optimization, particularly in feed-forward networks with ReLU activations. Orthogonal initialization (Saxe, McClelland, and Ganguli 2014), which ensures that all singular values of weights are 1, achieves dynamical isometry, the ideal state for trainability. In classifier pruning, TPP (Wang and Fu 2023) recently highlighted the importance of preserving trainability in pruned networks to support effective performance recovery.
@@
Previous works in the field of generative model compression (Chung et al. 2024; Fang, Ma, and Wang 2023) have shown that well-designed pruning techniques can lead compressed models to better solutions compared to their randomly initialized counterparts, while reaching the same performance more quickly (×$\times 2.5$ for StyleGAN2, ×$\times 8.0$ for DDPM). This is because pruned weights provide the compressed model with an initialization state that preserves the pre-trained model’s performance, starting with better performance compared to the random counterpart. This initial gain continues throughout the fine-tuning process, ultimately resulting in a better-performing compressed model. However, we observe that as fine-tuning progresses, models initialized with pruned weights converge more slowly compared to those initialized randomly (see Figure 2: orange line (DCP-GAN) vs. blue line (StyleKD)). This issue worsens as the model size decreases, resulting in suboptimal solutions, which indicates that the pruned weights themselves contain factors that hinder fine-tuning efficiency.
@@
We analyze the learned prior of pruned weights inherited from the pre-trained model by employing Singular Value Decomposition (SVD), which is widely adopted for low-rank approximation by identifying important basis from the weights (Denton et al. 2014; Zhang et al. 2015; Girshick 2015; Yoo et al. 2019). For a weight matrix $W\in\mathcal{R}^{m\times n}\ (m<n)$, we can decompose it using SVD:
@@
where $W^{\dagger}\in\mathcal{R}^{n\times m}$ is the Moore-Penrose inverse (Petersen 2012) and $\Sigma^{-1}=diag(\sigma^{+}_{1},\sigma^{+}_{2},\cdots,\sigma^{+}_{m})\in\mathcal{R}^{n\times m}$ with
@@
GS (Wang et al. 2020) 4.9M 1.3B 10.23 0.702 0.430 0.845 0.721
CAGAN (Liu et al. 2021) 9.23 0.685 0.500 0.760 0.722
StyleKD (Xu et al. 2022) 7.01 0.707 0.543 0.900 0.783
DCP-GAN (Chung et al. 2024) 8.17 0.680 0.539 0.787 0.741
@@
For the StyleGAN2 compression, we re-implement previous StyleGAN2 compression methods, which were implemented on the unofficial StyleGAN2 implementation [^1] https://github.com/rosinality/stylegan2-pytorch, on the official StyleGAN2 implementation [^2] https://github.com/NVlabs/stylegan2-ada-pytorch because the official implementation provides more optimized StyleGAN2 training settings and are easy to use. we apply our method to the weights pruned by DCP-GAN (Chung et al. 2024), which is the state-of-the-art pruning method in StyleGAN2 compression. We mainly compare our method against StyleKD (Xu et al. 2022) and DCP-GAN because they are the methods that demonstrate the best performance in StyleGAN2 compression. We use two StyleGAN2 architectures: the optimized architecture provided by the official implementation (denoted as “StyleGAN2 (small)”) and the original structure used in previous works (denoted as “StyleGAN2 (base)”). Experiments are conducted on the FFHQ (Karras, Laine, and Aila 2019) and LSUN Church (Yu et al. 2015) datasets. Following DCP-GAN, we use a facial mask as the content mask for the FFHQ dataset and a uniform mask (all pixel values are assigned values 1) for LSUN Church dataset. For StyleGAN3 compression, we implement our method based on the DCP-GAN implementation [^3] https://github.com/jiwoogit/DCP-GAN/tree/stylegan3. Similar to StyleGAN2 compression, we apply our method to the pruned weights using the DCP-GAN method for our experiments. We follow the experimental setup of DCP-GAN and train StyleGAN3 on the FFHQ dataset until the discriminator see 10 million images for both DCP-GAN and our method. For DDPM compression, we build upon Diff-Prune (Fang, Ma, and Wang 2023) [^4] https://github.com/VainF/Diff-Pruning, a foundational work in Diffusion model pruning. We train pruned DDPMs using the CIFAR10 (Alex 2009) and CelebA-HQ (Liu et al. 2015) datasets, following the experimental setup of Diff-Prune. Unlike Diff-Prune, which tested only on a 30% channel sparsity, we explore more extreme levels of sparsity in our experiments, including 50% and 70% channel sparsity. We provide more detailed implementation in the supplementary material.
@@
For StyleGAN compression, we evaluate our method using the Fréchet Inception Distance (FID) (Heusel et al. 2017) and Precision and Recall (P&R) (Kynkäänniemi et al. 2019). We also measure Density and Coverage (D&C) (Naeem et al. 2020), which are more robust to outliers than Precision and Recall. Here, Precision and Density evaluate the fidelity of generated samples, while Recall and Coverage assess their diversity. To calculate FID, we use all real samples from each dataset and 50K fake samples. For P&R and D&C calculation, we use 50K fake samples and all real samples from FFHQ, and 50K real samples from LSUN Church due to computational costs. For DDPM compression, in addition to FID, we employ Structural Similarity (SSIM) (Wang et al. 2004) to evaluate consistency with the pre-trained model, following the Diff-Prune. The SSIM score is measured between images generated by the pre-trained model and the compressed model, given the identical noise inputs. We use a 100-step DDIM sampler (Song, Meng, and Ermon 2021) for sampling.
@@
There are only a few pruning methods for generative models, with CAGAN (Liu et al. 2021), DCP-GAN, and Diff-Prune considered key baselines. We extend our analysis to CAGAN to further validate the applicability of SVS beyond state-of-the-art methods. The pruned weights of CAGAN exhibit a similar singular value distribution to DCP-GAN, with existence of dominant singular vectors. As shown in Table 4, applying SVS to CAGAN consistently leads to substantial improvements in all metrics, confirming that applicability of our method.
@@
In Table 6, we explore additional approaches to refine singular values: (a) normalizing singular values (setting $\sigma_{i}=1$), (b) normalizing with the spectral norm (dividing all singular values $\sigma_{i}$ by the largest singular value $\sigma_{1}$), (c) Squaring singular values, (d) applying spectral normalization to the synthesis network (Miyato et al. 2018), and (e) iterative refining the generator’s weights with our method. For (e), we perform iterative refining only for the first half of the training due to performance degradation. From (a) and (b), we find that setting singular values to 1 (where the weights are orthogonal) or spectral normalizing singular values cause the training to diverge, similar to previous observations (Saxe, McClelland, and Ganguli 2014; Wang and Fu 2023). From (c), we find that widening singular value gaps leads to early training divergence, highlighting the importance of minimizing these gaps for effective fine-tuning. From (d), we see that applying spectral normalization to the generator’s weights makes it difficult for the generator to learn. Finally, from (e), we find that repeatedly refining the singular values of weights during training perturbs the learning process, depriving the generator of sufficient training time and leading to suboptimal results. From these results, we find that applying our method once at model initialization is not only efficient but also leads to more effective training.
@@ Conclusion
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2020-0-01336, Artificial Intelligence Graduate School Program(UNIST)), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT). (No.2022R1C1C100849612), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2022-0-00959, (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2022-II220264, Comprehensive Video Understanding and Generation with Knowledge-based Deep Logic Neural Network) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-02068, Artificial Intelligence Innovation Hub).

### v2

@@ Abstract
While pruning methods effectively maintain model performance without extra training costs, they often focus solely on preserving crucial connections, overlooking the impact of pruned weights on subsequent fine-tuning or distillation, leading to inefficiencies. Moreover, most compression techniques for generative models have been developed primarily for GANs, tailored to specific architectures like StyleGAN, and research into compressing Diffusion models has just begun. Even more, these methods are often applicable only to GANs or Diffusion models, highlighting the need for approaches that work across both model types. In this paper, we introduce Singular Value Scaling (SVS), a versatile technique for refining pruned weights, applicable to both model types. Our analysis reveals that pruned weights often exhibit dominant singular vectors, hindering fine-tuning efficiency and leading to suboptimal performance compared to random initialization. Our method enhances weight initialization by minimizing the disparities between singular values of pruned weights, thereby improving the fine-tuning process. This approach not only guides the compressed model toward superior solutions but also significantly speeds up fine-tuning. Extensive experiments on StyleGAN2, StyleGAN3 and DDPM demonstrate that SVS improves compression performance across model types without additional training costs. Our code is available at: https://github.com/LAIT-CVLab/Singular-Value-Scaling.
@@ Introduction
Generative models like Generative Adversarial Networks (GANs) [Goodfellow et al. 2014] and Diffusion models [Ho et al. 2020] have achieved remarkable performance across various computer vision tasks such as image generation [Karras et al. 2020, Ho et al. 2020, Wang et al. 2022], image editing [Baykal et al. 2023, Pehlivan et al. 2023, Kawar et al. 2023, Zhang et al. 2023], even video generation [Skorokhodov et al. 2022, Ho et al. 2022, Blattmann et al. 2023] and 3D generation [Chan et al. 2022, Karnewar et al. 2023]. The impressive performance of these generative models, however, often comes at the cost of high memory and computational demands, limiting their real-world applicability. To address these issues, several model compression technique for generative models have been proposed [Liu et al. 2021, Xu et al. 2022, Chung et al. 2024, Fang et al. 2023].
Model compression typically involves two steps: 1) pruning to reduce model size while retaining essential pre-trained knowledge, and 2) fine-tuning to restore performance. Among these, effective pruning has received significant attention for its ability to enhance following fine-tuning step. The retained pre-trained knowledge from pruning can enhance fine-tuning, leading to improved performance and faster convergence without additional training costs. While previous pruning methods [Chung et al. 2024, Fang et al. 2023] effectively maintain model performance, they often focus solely on preserving crucial connections in the pre-trained model, overlooking the impact of pruned weights on subsequent process, leading to inefficient fine-tuning or distillation. This issue becomes more severe as the model’s capacity decreases, and in some cases, the pruned weights results in worse performance compared to random initialization; i.e. slow convergence speed and lower performance. Therefore, addressing these factors is essential for achieving more efficient generative models.
@@
Recently, several studies [Liu et al. 2021, Xu et al. 2022, Chung et al. 2024] have been proposed for compressing unconditional GANs, particularly the StyleGAN family [Karras et al. 2020, 2021], which are the state-of-the-art models in the area. CAGAN [Liu et al. 2021] introduced the first framework for pruning pre-trained StyleGAN models and fine-tuning them via pixel-level and feature-level knowledge distillation. Following this, StyleKD [Xu et al. 2022] further enhanced the fine-tuning process by incorporating a specialized relation loss tailored for StyleGAN. In particular, CAGAN introduced a content-aware pruning technique that preserves connections in the pre-trained model crucial for the semantic part of generated images, while StyleKD tackled output discrepancies problem by inheriting only the pre-trained model’s mapping network and randomly initializing the synthesis network with smaller size. More recently, DCP-GAN [Chung et al. 2024] proposed a diversity-aware channel pruning technique, which preserves connections in the synthesis network that contribute to sample diversity, significantly enhancing both the diversity of generated images and the training speed.
@@
Diffusion models have garnered significant attention for their stable training and impressive generative capabilities. However, their performance incurs high computational costs due to iterative sampling. To enhance efficiency, various sampling techniques have been developed to reduce the number of required iterations [Song et al. 2021, Lu et al. 2022, Zheng et al. 2023]. Orthogonal to the these sampling techniques, Diff-Prune [Fang et al. 2023] introduced a seminal work for reducing computational costs by compressing Diffusion models. This demonstrated that by primarily pruning connections involved in less important diffusion steps, a reduced-size Diffusion model with minimal performance loss can be achieved in significantly fewer training iterations compared to the original-sized model.
@@
Weight initialization is essential for efficient training in deep learning. Glorot [Glorot and Bengio 2010] and He initialization [He et al. 2015] are widely used to maintain activation variance and ensure stable optimization, particularly in feed-forward networks with ReLU activations. Orthogonal initialization [Saxe et al. 2014], which ensures that all singular values of weights are 1, achieves dynamical isometry, the ideal state for trainability. In classifier pruning, TPP [Wang and Fu 2023] recently highlighted the importance of preserving trainability in pruned networks to support effective performance recovery.
@@
Previous works in the field of generative model compression [Chung et al. 2024, Fang et al. 2023] have shown that well-designed pruning techniques can lead compressed models to better solutions compared to their randomly initialized counterparts, while reaching the same performance more quickly (×$\times 2.5$ for StyleGAN2, ×$\times 8.0$ for DDPM). This is because pruned weights provide the compressed model with an initialization state that preserves the pre-trained model’s performance, starting with better performance compared to the random counterpart. This initial gain continues throughout the fine-tuning process, ultimately resulting in a better-performing compressed model. However, we observe that as fine-tuning progresses, models initialized with pruned weights converge more slowly compared to those initialized randomly (see Figure 2: orange line (DCP-GAN) vs. blue line (StyleKD)). This issue worsens as the model size decreases, resulting in suboptimal solutions, which indicates that the pruned weights themselves contain factors that hinder fine-tuning efficiency.
@@
We analyze the learned prior of pruned weights inherited from the pre-trained model by employing Singular Value Decomposition (SVD), which is widely adopted for low-rank approximation by identifying important basis from the weights [Denton et al. 2014, Zhang et al. 2015, Girshick 2015, Yoo et al. 2019]. For a weight matrix $W\in\mathcal{R}^{m\times n}\ (m<n)$, we can decompose it using SVD:
@@
where $W^{\dagger}\in\mathcal{R}^{n\times m}$ is the Moore-Penrose inverse [Petersen 2012] and $\Sigma^{-1}=diag(\sigma^{+}_{1},\sigma^{+}_{2},\cdots,\sigma^{+}_{m})\in\mathcal{R}^{n\times m}$ with
@@
GS [Wang et al. 2020] 4.9M 1.3B 10.23 0.702 0.430 0.845 0.721
CAGAN [Liu et al. 2021] 9.23 0.685 0.500 0.760 0.722
StyleKD [Xu et al. 2022] 7.01 0.707 0.543 0.900 0.783
DCP-GAN [Chung et al. 2024] 8.17 0.680 0.539 0.787 0.741
@@
For the StyleGAN2 compression, we re-implement previous StyleGAN2 compression methods, which were implemented on the unofficial StyleGAN2 implementation [^1] https://github.com/rosinality/stylegan2-pytorch, on the official StyleGAN2 implementation [^2] https://github.com/NVlabs/stylegan2-ada-pytorch because the official implementation provides more optimized StyleGAN2 training settings and are easy to use. we apply our method to the weights pruned by DCP-GAN [Chung et al. 2024], which is the state-of-the-art pruning method in StyleGAN2 compression. We mainly compare our method against StyleKD [Xu et al. 2022] and DCP-GAN because they are the methods that demonstrate the best performance in StyleGAN2 compression. We use two StyleGAN2 architectures: the optimized architecture provided by the official implementation (denoted as “StyleGAN2 (small)”) and the original structure used in previous works (denoted as “StyleGAN2 (base)”). Experiments are conducted on the FFHQ [Karras et al. 2019] and LSUN Church [Yu et al. 2015] datasets. Following DCP-GAN, we use a facial mask as the content mask for the FFHQ dataset and a uniform mask (all pixel values are assigned values 1) for LSUN Church dataset. For StyleGAN3 compression, we implement our method based on the DCP-GAN implementation [^3] https://github.com/jiwoogit/DCP-GAN/tree/stylegan3. Similar to StyleGAN2 compression, we apply our method to the pruned weights using the DCP-GAN method for our experiments. We follow the experimental setup of DCP-GAN and train StyleGAN3 on the FFHQ dataset until the discriminator see 10 million images for both DCP-GAN and our method. For DDPM compression, we build upon Diff-Prune [Fang et al. 2023] [^4] https://github.com/VainF/Diff-Pruning, a foundational work in Diffusion model pruning. We train pruned DDPMs using the CIFAR10 [Alex 2009] and CelebA-HQ [Liu et al. 2015] datasets, following the experimental setup of Diff-Prune. Unlike Diff-Prune, which tested only on a 30% channel sparsity, we explore more extreme levels of sparsity in our experiments, including 50% and 70% channel sparsity. We provide more detailed implementation in the supplementary material.
@@
For StyleGAN compression, we evaluate our method using the Fréchet Inception Distance (FID) [Heusel et al. 2017] and Precision and Recall (P&R) [Kynkäänniemi et al. 2019]. We also measure Density and Coverage (D&C) [Naeem et al. 2020], which are more robust to outliers than Precision and Recall. Here, Precision and Density evaluate the fidelity of generated samples, while Recall and Coverage assess their diversity. To calculate FID, we use all real samples from each dataset and 50K fake samples. For P&R and D&C calculation, we use 50K fake samples and all real samples from FFHQ, and 50K real samples from LSUN Church due to computational costs. For DDPM compression, in addition to FID, we employ Structural Similarity (SSIM) [Wang et al. 2004] to evaluate consistency with the pre-trained model, following the Diff-Prune. The SSIM score is measured between images generated by the pre-trained model and the compressed model, given the identical noise inputs. We use a 100-step DDIM sampler [Song et al. 2021] for sampling.
@@
There are only a few pruning methods for generative models, with CAGAN [Liu et al. 2021], DCP-GAN, and Diff-Prune considered key baselines. We extend our analysis to CAGAN to further validate the applicability of SVS beyond state-of-the-art methods. The pruned weights of CAGAN exhibit a similar singular value distribution to DCP-GAN, with existence of dominant singular vectors. As shown in Table 4, applying SVS to CAGAN consistently leads to substantial improvements in all metrics, confirming that applicability of our method.
@@
In Table 6, we explore additional approaches to refine singular values: (a) normalizing singular values (setting $\sigma_{i}=1$), (b) normalizing with the spectral norm (dividing all singular values $\sigma_{i}$ by the largest singular value $\sigma_{1}$), (c) Squaring singular values, (d) applying spectral normalization to the synthesis network [Miyato et al. 2018], and (e) iterative refining the generator’s weights with our method. For (e), we perform iterative refining only for the first half of the training due to performance degradation. From (a) and (b), we find that setting singular values to 1 (where the weights are orthogonal) or spectral normalizing singular values cause the training to diverge, similar to previous observations [Saxe et al. 2014, Wang and Fu 2023]. From (c), we find that widening singular value gaps leads to early training divergence, highlighting the importance of minimizing these gaps for effective fine-tuning. From (d), we see that applying spectral normalization to the generator’s weights makes it difficult for the generator to learn. Finally, from (e), we find that repeatedly refining the singular values of weights during training perturbs the learning process, depriving the generator of sufficient training time and leading to suboptimal results. From these results, we find that applying our method once at model initialization is not only efficient but also leads to more effective training.